Two cars, A and B, are held by the two pistons of a hydraulic press at the same

horizontal level. Car A weighs 3,400 N, and car B weighs 2,850 N. Calculate the

ratio between the area of the piston car A is sitting on and the area of the piston

car B is sitting on.


1
Expert's answer
2022-01-09T18:23:57-0500

"F_A=3400N,F_B=2850N"

Area of car A

"A_A"

Area of car B

"A_B"

Pascal law

"P_A=P_B"

"\frac{F_A}{A_A}=\frac{F_B}{A_B}"

"\frac{A_A}{A_B}=\frac{F_A}{F_B}"

"\frac{A_A}{A_B}=\frac{3400}{2850}=1.192"

Ratio of areas

"\frac{A_A}{A_B}=1.192"